Suppose that type I sellers charged the price of $60 for the portable TV, type II sellers charged $80, type III sellers charged $100, type IV sellers charged $120, and type V sellers charged $140. 


Determine 

the expected lowest price for the TV from one, two, three, four, and five searches and

the marginal benefit from each additional search.


1
Expert's answer
2021-07-02T11:39:38-0400

A) Type 1= $60 type 2= $80 type 3= $100 4th =$120 5th=$140

find the RANGE:

The range = the highest – the lowest "=140 – 60= \$80"

lowest expected price at 1st search "= 60 +\frac{ range }{ (number \space of\space searches +1)}"

Lowest expected price at the 1st search "= 60 + \frac{80 }{1+1} = 100"

2nd search "=60+\frac{ 80}{2+1} = 86.67"

3rd search "= 60+ \frac{8}{3+1}= 80"

4th search "=60+\frac{8}{4+1} = 76"

5th search "= 60+\frac{8}{5+1}= 73.33"


B) Marginal benefit = 1st expected lowest price – 2nd expected lowest price

We start from second search because first one always equals = 0

Second search"=100-86.67=\$13.33"

Third search"= 86.67 – 80 = \$6.67"

Fourth search"= 80 – 76 = \$4"

Fifth search "= 76 – 73.33 = \$2.67"